OpenStudy (maeganr):

What is an equation in slope-intercept form on the line that passes through (6,10) and is parallel to the graph of y = 1/3x - 1?

OpenStudy (anonymous):

take co ordinate geometry in use

OpenStudy (anonymous):

Do you know what slope intercept form is? Do you know what the slopes of parallel lines have in common?

OpenStudy (maeganr):

no not really

OpenStudy (anonymous):

10 = (1/3)(6) + b 10 = 2 + b b = 8 so y = 1/3x + 8 is answer

OpenStudy (anonymous):

Well @Argos gave you the answer. If you need further explanation, I can explain.

OpenStudy (maeganr):

Lil bit. Im in 7th grade and doing 9th grade work and i dont get this crap. ):

OpenStudy (anonymous):

can u medal me pls if i gave u good answer otherwise my mom will make me get off computer im only 10

OpenStudy (anonymous):

Parallel lines have the same slope y= 1/3x - 1 Slope intercept form: y=mx+b, and m is the slope For the equation above, whats the slope?

OpenStudy (maeganr):

1/3

OpenStudy (anonymous):

The slopes of parallel lines have the same slope. The problem tells you that the line you find must be parallel to y =1/3x - 1..therefore the slope of the line you find must have the same slope as y = 1/3x - 1

OpenStudy (anonymous):

Do you know point slope form?

OpenStudy (maeganr):

i dont know what i know

OpenStudy (maeganr):

im so confused

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Does this look familiar? \[ y- y _{1}= m ( x - x _{1})\]

OpenStudy (maeganr):

Yes, but we do it differently, we divide

OpenStudy (anonymous):

divide for point slope form, or for finding slope?

OpenStudy (maeganr):

slope formula

OpenStudy (anonymous):

No, this is something else called point slope formula, when you need to use a point and slope to find the equation of a line

OpenStudy (maeganr):

Ok I think I am getting it now, I was absent and missed the notes and lesson on this

OpenStudy (anonymous):

Ok, just let me know if you dont understand anything. So can you tell me what point you need to use, and what slope?

OpenStudy (maeganr):

Maybe I do not understand, The only information I had is what I typed initially. I am sorry, I won't take up your time

OpenStudy (anonymous):

You gave the point: (6,10) And said the parallel slope was 1/3, right?

OpenStudy (maeganr):

yes

OpenStudy (anonymous):

Do you know what y1 and x1 are?

OpenStudy (maeganr):

yes, but I do not know how to get them for this equation

OpenStudy (anonymous):

y1 is just the y part of the point. You do know that a point is (x,y)?

OpenStudy (maeganr):

yes

OpenStudy (maeganr):

So I plug 6 and 10 into the x & y and complete the equation?

OpenStudy (anonymous):

Yeah, and plug in the 1/3 for the slope (m)

OpenStudy (maeganr):

ok

OpenStudy (anonymous):

Let me know what you get

OpenStudy (maeganr):

10 = 1?

OpenStudy (anonymous):

o_o can you tell me how the equation looked once it was all plugged in?

OpenStudy (maeganr):

\[10 = 1/3 (6)- 1\]

OpenStudy (anonymous):

What happened to the y and the x?

OpenStudy (maeganr):

10 = 2-1 10/1 = 1/1 10 =1

OpenStudy (maeganr):

I replaced them with the points (6, 10)

OpenStudy (anonymous):

No.. wait let me show you with an example: Get it?